a pharmacist wishes to mix a solution that is 3% minoxidil she has on hand 60 ML of a 2 % solution she wishes to add some 5% solution to obtain the desired 3%solution. how much 5%solution should she add?
Answer by lwsshak3(11628)

a pharmacist wishes to mix a solution that is 3% minoxidil she has on hand 60 ML of a 2 % solution she wishes to add some 5% solution to obtain the desired 3%solution. how much 5%solution should she add?
**
let x=ml of 5% to be added
60+x=ml of the desiredl 3% solution
..
2%60+5%x=3%(60+x)
.02*60+.05x=.03*60+.03x
1.2+.05x=1.8+.03x
.02x=.6
x=.6/.02=30 ml
ans:
30 ml of the 5% solution must be added to 60 ml of the 2% solution obtain the desired 3% solution.